The sides of two regular octagons are 3 inches and 8 inches. Find the ratio of

(a) their perimeters, (b) their radii, and (c) their areas.



1
Expert's answer
2021-01-27T01:16:10-0500

Solution. a) Using perimeter definition


"\frac{P_2}{P_1}=\frac{8a_2}{8a_1}=\frac{a_2}{a_1}=\frac{8}{3}"

where a2=8 inches and a1=3 inches are sides of two regular octagons.

b) The radii of the inscribed and circumscribed circles are related to the side of a regular octagon by the formulas


"R=\frac{a}{2sin22.5^0}"

"r=\frac{a}{2tan22.5^0}"

where R and r are radii of regular octagon; a is side of the regular octagon. Therefore, the ratio of the radii is proportional to the ratio of the sides and is equal to


"\frac{R_2}{R_1}=\frac{r_2}{r_1}=\frac{a_2}{a_1}=\frac{8}{3}"

с) We use the formula for the area of a regular octagon


"A=2a^2(1+\sqrt{2})"

where A is area of a regular octagon; a is side of the regular octagon. Therefore


"\frac{A_2}{A_1}=\frac{2a^2_2(1+\sqrt{2})}{2a^2_1(1+\sqrt{2})}=\frac{a^2_2}{a^2_1}=\frac{64}{9}"

Answer. a)


"\frac{P_2}{P_1}=\frac{8}{3}"

b)


"\frac{R_2}{R_1}=\frac{r_2}{r_1}=\frac{a_2}{a_1}=\frac{8}{3}"

c)

"\frac{A_2}{A_1}=\frac{64}{9}"